# Reweighted Spectral Partitioning Works: Bounds for Special Graph Classes

### v3

@@ 3.2 Bounded-Genus Graphs
where the antepenultimate inequality follows from Cauchy-Schwarz, and the penultimate inequality follows from the triangle inequality. ∎
Lemma 3.12.
@@
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices and maximum degree 444 with genus $g$ so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
@@
[…]
@@
Figure 3: Two examples of the degree-reduction construction for a $g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in some degree-111 vertices).
@@
We will be able to leverage this lemma in order to turn a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of constant maximum degree into a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. In particular, we will obtain a bound on $\lambda_{2}^{*}$ (and $\gamma^{(1)}$) for triangulated bounded-genus graphs of constant degree by adapting Kelner’s techniques [Kel04, Kel06b, Kel06a] for bounding $\lambda_{2}$. Before exploring these techniques, we first show how to reduce to the case of triangulated bounded-genus graphs, which will use a similar technique:
Lemma 3.13.
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated $g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $O(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Figure 4: A demonstration of the triangulation steps on a face. From left to right, top to bottom: $G$, $H_{1}$, $H_{2}$, $H_{3}$, $H$.
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a $g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth 222, and hence Lemma 3.11 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size >4>4> 4. Moreover, any vertex incident to a face of size >4>4> 4 (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly 333. For each face of exactly size 444 in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 4 for examples of each of these graphs around a particular face. It’s worth noting that for non-cellular embeddings of higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). Fortunately, this does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most 222 edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated $g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 4 for a demonstration of this algorithm.
@@
For the remaining lemmas, we need only consider triangulated $g$ graphs.
A $g$ graph naturally embeds into a $g$ surface. Moreover, a $g$ graph in fact has a circle packing on a $g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Theorem 3.5, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most 333) instead of $\lambda_{2}$.
The subdivision procedure used by Kelner is the hexagonal subdivision of a triangulation $G$, which produces a graph $G^{\prime}$ that by bisecting every edge in $E(G)$, and then connecting all three bisection vertices around each resulting face to form a triangle with new edges. See Figure 5 for an example of this operation.
@@
We defer the proof of Lemma 3.14 to Appendix C since it is quite similar to that of Lemma 3.11, and will primarily make use of an argument of Kelner  [Kel06a, Kel04, Kel06b] that randomly “embeds” $G$ into $G^{(k)}$. Compared to the proof of Lemma 3.11, this randomized argument includes random paths in addition to random vertices.
@@
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a $g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
@@
Apply Lemma 3.15 with $\varepsilon=O(1)$ to get $N$. Fix any $k\geq N$. Denote $n^{(k)}:=|V(G^{(k)})|$. We will now choose the values $y,f$ for the formulation of $\dot{\gamma}^{(3)}$ to obtain a similar argument to the proof of Theorem 3.5. For each $v\in V(G^{(k)})$, choose $s(v):=D(f^{(k)}(C_{v}))$ for some constant $\rho\in\Theta(1)$ to be chosen later. Choose $f(v)$ as the image of the centre of $C_{v}$ under $f^{(k)}$. Note that under this choice, $\sum_{v\in V(G^{(k)})}f^{(k)}(v)$ is exactly the origin.
@@
The combination of Lemma 3.12, Lemma 3.13, Lemma 3.14 and Lemma 3.16 together imply the following theorem:
Theorem 3.17.
@@
$\frac{\lambda_{2}(G)}{\Delta}\leq\lambda_{2}^{*}(G)\leq\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$

### v4

@@ 5.3.1 Uniform Shallow Minors
where the preantepenultimate step follows from Cauchy-Schwarz, the antepenultimate step follows from the triangle inequality, and the last step uses $k\leq 2L+1$. Thus, the edge constraints are satisfied.
It remains to verify the objective. To obtain a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$, we first bound the squared distances within each $p^{-1}(v)$. For any vertex $v\in V$ and any two vertices $u^{\prime},v^{\prime}\in p^{-1}(v)$, there is a path $u^{\prime}=v_{1}^{\prime},\dots,v_{l+1}^{\prime}=v^{\prime}$ of length $l\leq L$ contained entirely within $p^{-1}(v)$. By the triangle inequality, Cauchy-Schwarz, and the constraint from $\gamma^{(1)}(H)$,
$\begin{array}[]{r c l}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}&\leq&\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|\right]^{2}\\[4.0pt] &\leq&l\left[|f_{H}(v_{1}^{\prime})-f_{H}(v_{2}^{\prime})|^{2}+\cdots+|f_{H}(v_{l}^{\prime})-f_{H}(v_{l+1}^{\prime})|^{2}\right]\\[4.0pt] &\leq&L\sum_{i=1}^{l}\left[y_{H}(v_{i}^{\prime})+y_{H}(v_{i+1}^{\prime})\right].\end{array}$
Summing over all pairs $u^{\prime},v^{\prime}\in p^{-1}(v)$ and noting that each $y_{H}(w^{\prime})$ for $w^{\prime}\in p^{-1}(v)$ appears in at most $2|p^{-1}(v)|^{2}$ terms,
$\sum_{u^{\prime}\in p^{-1}(v)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\leq 2Lr^{2}\sum_{w^{\prime}\in p^{-1}(v)}y_{H}(w^{\prime}).$
Using this bound, we obtain a lower bound on $\sum_{v\in V}|f_{G}(v)|^{2}$:
$\begin{array}[]{r c l}2|V|\cdot\sum_{v\in V}|f_{G}(v)|^{2}&=&\sum_{u,v\in V}|f_{G}(u)-f_{G}(v)|^{2}\\[12.0pt] &=&\sum_{u,v\in V}|f_{H}(m(u))-f_{H}(m(v))|^{2}\\[12.0pt] &\geq&\operatorname{\mathbb{E}}\left[\sum_{u,v\in V}|f_{H}(\pi(u))-f_{H}(\pi(v))|^{2}\right]\\[12.0pt] &=&\frac{1}{r^{2}}\sum_{u\neq v\in V}\sum_{u^{\prime}\in p^{-1}(u)}\sum_{v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\\[12.0pt] &=&\frac{1}{r^{2}}\left[\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}-\sum_{v\in V}\sum_{u^{\prime},v^{\prime}\in p^{-1}(v)}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}\right]\\[12.0pt] &\geq&\frac{1}{r^{2}}\left[2r|V|\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}-2Lr^{2}\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})\right]\\[12.0pt] &=&\frac{2\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}\left[|V|-Lr\gamma^{(1)}(H)\right]\\[12.0pt] &\geq&\frac{2(|V|-1)\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r},\end{array}$
where the antepenultimate step uses $\sum_{u^{\prime},v^{\prime}\in V^{\prime}}|f_{H}(u^{\prime})-f_{H}(v^{\prime})|^{2}=2|V^{\prime}|\sum|f_{H}(v^{\prime})|^{2}$ and the bound on within-set distances, the penultimate step uses $\sum y_{H}(v^{\prime})=\gamma^{(1)}(H)\sum|f_{H}(v^{\prime})|^{2}$, and the last step uses the assumption $Lr\gamma^{(1)}(H)\leq 1$ from the beginning of the proof. Thus, $\sum_{v\in V}|f_{G}(v)|^{2}\gtrsim\frac{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}}{r}$.
Finally, we verify the objective:
$\begin{array}[]{r c l}\frac{\sum_{v\in V}y_{G}(v)}{\sum_{v\in V}|f_{G}(v)|^{2}}&=&\frac{2(2L+1)\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v\in V}|f_{G}(v)|^{2}}\\[12.0pt] &\lesssim&\frac{2(2L+1)\sum_{v^{\prime}\in V^{\prime}}y_{H}(v^{\prime})}{\sum_{v^{\prime}\in V^{\prime}}|f_{H}(v^{\prime})|^{2}/r}\\[12.0pt] &=&2r(2L+1)\cdot\gamma^{(1)}(H)\\[12.0pt] &\lesssim&r\cdot L\cdot\gamma^{(1)}(H).\end{array}$
We will now apply this result to prove the following lemma:
Lemma 5.15 (Restatement of Lemma 2.8).
@@
Let $G$ be a graph with genus $g$, $n$ vertices, and maximum degree $\Delta$. Then there exists a graph $H$ with $n\Delta$ vertices, maximum degree $4$, and genus $g$, so that $\gamma^{(1)}(G)\lesssim\gamma^{(1)}(H)\cdot\Delta\cdot\log\Delta$.
@@
We create a set $V^{\prime}$ that contains $\Delta$ vertices for each vertex in $V(G)$. Let $p:V^{\prime}\to V(G)$ denote the mapping of these vertices. Likewise, for each edge-vertex pair $(e,v)\in E\times V$ with $e$ incident to $v$, map $(e,v)$ to some vertex in $p^{-1}(v)$, in such a manner that no two distinct edge-vertex pairs are mapped to the same vertex in $V^{\prime}$. Use this second mapping to construct a set of endpoint-distinct edges $E^{\prime}$. For each $v\in V(G)$, construct a near-perfect binary tree $T_{v}$ over the $\Delta$ vertices in $p^{-1}(v)$ so that some Euler tour of $T_{v}$ contains the vertices mapped to the rotation system around $v$ as a subsequence (it suffices to construct a near-perfect binary search tree over an arbitrary ordering of $p^{-1}(v)$ containing this subsequence). Note that $T_{v}$ will have max-degree $3$ and diameter at most $2\log_{2}\Delta$. Let $E^{\prime\prime}$ be the union over all such trees. Let $H:=(V^{\prime},E^{\prime}\cup E^{\prime\prime})$. Note that the maximum degree of $H$ is $4$, since each vertex in $V^{\prime}$ is incident to at most one edge in $E^{\prime}$ and at most $3$ edges in $E^{\prime\prime}$.
@@
Figure 8: Two examples of the degree-reduction construction for a genus-$g$ graph, respecting a rotation system, local to a vertex. The vertex is replaced with $\Delta$ vertices, at least one vertex for each incident edge, and these new vertices are connected with a near-perfect binary tree. In the left example, the displayed vertex is of maximum degree in its graph, while in the right example it is not (resulting in a degree-11 vertex).
@@
We will be able to leverage this lemma in order to turn a bound on $\gamma^{(1)}$ for bounded-genus graphs of constant maximum degree into a bound on $\gamma^{(1)}$ for bounded-genus graphs of arbitrary maximum degree at only a $\log\Delta$ loss. We will now use a similar technique to further reduce to the case of triangulated bounded-genus graphs with constant maximum degree:
Lemma 5.16 (Restatement of Lemma 2.9).
Let $G$ be a graph of genus $g$ with $n$ vertices and maximum degree $\Delta$. Then there is a triangulated genus-$g$ graph $H$ with $(\Delta+1)\cdot n$ vertices and maximum degree $\operatorname{\mathcal{O}}(\Delta)$, so that $\gamma^{(1)}(G)\lesssim(\Delta+1)\cdot\gamma^{(1)}(H)$.
@@
Figure 9: A demonstration of the triangulation steps on a face. From left to right, top to bottom: $G$, $H_{1}$, $H_{2}$, $_{3}, HH$. The same steps can also be performed on a non-cellular face.
Let $H_{1}$ be the graph that adds $\Delta$ vertices around each vertex to $G$, and ensure that between any two edges in the rotation system around a vertex in $G$, at least one such edge is added. Observe that $H_{1}$ is a genus-$g$ graph with $(\Delta+1)n$ vertices and maximum degree $2\Delta$. Observe also that $G$ is a uniform shallow minor of $H_{1}$ of depth $2$, and hence Lemma 5.14 shows that $\gamma^{(1)}(G)\lesssim(\Delta+1)\gamma^{(1)}(H_{1})$.
Now, we will add edges to $H_{1}$. In particular, we will construct a sequence of graphs that add edges to $H_{1}$, but never add vertices, and hence $\gamma^{(1)}$ will only increase for this sequence of graphs. Specifically, within each face of $H_{1}$, we create a cycle of all the newly added vertices in order around the face. The result is a graph we call $H_{2}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{3,2\Delta\}$) that has no adjacent pair of faces both with size >4>4. Moreover, any vertex incident to a face of size >4>4 (i.e., those in the created cycles) is incident exactly two other faces, and has degree exactly $3$. For each face of exactly size $4$ in $H_{2}$, triangulate it arbitrarily to get another new graph $H_{3}$ (with genus $g$, $(\Delta+1)n$ vertices, and maximum degree $\max\{5,4\Delta\}$). See Figure 9 for examples of each of these graphs around a particular face. It’s worth noting that for non-cellular embeddings of higher genus graphs, it is possible that a non-simple face may include the same edge twice in the same “direction” during a traversal around the face (as opposed to the example in the figure, where the doubly-included edge is included in opposite directions). Fortunately, this does not cause issues for the construction described here either. It is also worth noting that these steps are necessary even if all the faces are simple: Specifically, we would like to handle the remaining non-triangular faces independently, and the separation ensures that we are not able to later produce parallel edges (see [KB97, Figure 6.3(b)] for a demonstration of the issues of adjacent simple faces).
It remains to triangulate the remaining non-triangular faces of $H_{3}$, none of which are adjacent, and all of which are simple. This can be accomplished by adding at most $2$ edges per vertex. In particular, Kant and Bodlaender [KB92, KB97] gave an algorithm called “zig-zagging” that adds edges to a single face using a sequence of ear-cuts, triangulating the face. This algorithm applies in the case of faces on any rotation system, not just a rotation system on a planar graph. Call the final graph $H$, which is a triangulated genus-$g$ graph with maximum degree $\max\{7,4\Delta\}$. See the last diagram in Figure 9 for a demonstration of this algorithm.
@@
5.3.2 Circle Packings for Bounded-Genus Graphs
Because of the lemmas in Section 5.3.1 we henceforth need only consider triangulated genus-$g$ graphs of constant degree. We will now begin generalizing the arguments of Kelner [Kel04, Kel06b, Kel06a] to obtain bounds on $\gamma^{(1)}(G)$ for this class.
A genus-$g$ graph naturally embeds into a genus-$g$ surface. Moreover, a genus-$g$ graph in fact has a circle packing on a genus-$g$ surface. Both the surface and packing is unique up to a certain set of transformations [HS93]. However, to bound $\gamma^{(3)}$ in a similar manner to Proposition 2.4, we would need a circle packing on the unit sphere $S^{2}$, likely of ply $O(g)$. Unfortunately, as Kelner [Kel04, Kel06b, Kel06a] pointed out, such circle packings need-not exist. Instead, in order to bound $\lambda_{2}$, Kelner made use of two steps: First, a specific subdivision procedure is applied to the graph, and it is shown that a bound for the graph obtained by a sequence of subdivision procedures also induces a bound on the initial graph. Second, by applying this procedure a sufficient number of times, it is shown that a continuous analogue of the circle packing theory can be used to obtain a bound. We will apply analogues of each of these steps for $\gamma^{(3)}$ (and $\gamma^{(1)}$, which differs only by a constant factor of at most $3$) instead of $\lambda_{2}$.
The subdivision procedure used by Kelner is the hexagonal subdivision of a triangulation $G$, which produces a graph $G^{\prime}$ by bisecting every edge in $E(G)$, and then connecting all three bisection vertices around each resulting face to form a triangle with new edges. See Figure 10 for an example of this operation.
@@ 5.3.2 Circle Packings for Bounded-Genus Graphs
We defer the proof of Lemma 5.17 to Appendix C since it is similar to that of Lemma 5.14, and the details are quite technical. Essentially, it will make use of an argument of Kelner [Kel06a, Kel04, Kel06b] that randomly “embeds” $G$ into $G^{(k)}$. The main distinguishing feature compared to the proof of Lemma 5.14 is that this randomized argument includes not just random vertices, but also random paths between vertices.
@@
Let $G$ be a graph of genus $g$. For each $k$, let $F^{(k)}$ denote a genus-$g$ surface on which $G^{(k)}$ admits a circle packing. For each vertex $v\in G^{(k)}$, let $C_{v}$ denote the disk in the packing corresponding to $v$, and let $r_{v}$ denote its radius. Let $A(C)$ and $D(C)$ denote the area (measure) and diameter (longest geodesic path between a point pair), respectively, of a compact and connected region $C$ either in $F^{(k)}$ or $S^{2}$. Then there is a sequence $f^{(k)}:F^{(k)}\to S^{2}$ of analytic maps so that, for any $\varepsilon>0$, there exists a threshold $N$ such that for any $k\geq N$ a partition of $S^{(k)}$ into $S^{(k)}_{1}\cup S^{(k)}_{2}$ exists with the following properties:
@@
Apply Lemma 5.18 with $\varepsilon=O(1)$ to get $N$. Fix any $k\geq N$. Denote $n^{(k)}:=|V(G^{(k)})|$. We will now choose the values $y,f$ for the formulation of $\dot{\gamma}^{(3)}$ to obtain a similar argument to the proof of Proposition 2.4. For each $v\in V(G^{(k)})$, choose $s(v):=D(f^{(k)}(C_{v}))$. Choose $f(v)$ as the image of the centre of $C_{v}$ under $f^{(k)}$. Note that under this choice, $\sum_{v\in V(G^{(k)})}f^{(k)}(v)$ is exactly the origin.
@@
5.3.3 Combining the Lemmas
The combination of Lemma 2.8, Lemma 2.9, Lemma 5.17 and Lemma 5.19 together imply the main result:
Proposition 5.20 (Restatement of Proposition 2.7).
@@ 5.3.3 Combining the Lemmas
$\gamma^{(1)}(G)\lesssim\frac{g\log\Delta}{n}.$
Proof.
Let $G^{\prime}$ be the genus-$g$ graph on $n\Delta$ vertices with maximum degree at most $4=O(1)$ produced by Lemma 2.8. Since the maximum degree of $G^{\prime}$ is constant, a triangulated genus-$g$ graph $G^{\prime\prime}$ with $O(n\Delta)$ vertices and constant maximum degree can be obtained by Lemma 2.9 so that $\gamma^{(1)}(G^{\prime})\lesssim\gamma^{(1)}(G^{\prime\prime})$. By Lemma 5.17, and Lemma 5.19 $\gamma^{(1)}(G^{\prime\prime})\lesssim\frac{g}{n\Delta}$, since the degree of $G^{\prime\prime}$ is constant. Hence,
$\gamma^{(1)}(G)\lesssim\gamma^{(1)}(G^{\prime})\cdot\Delta\log\Delta\leq\gamma^{(1)}(G^{\prime\prime})\cdot\Delta\log\Delta\lesssim\frac{g\log\Delta}{n}.$
6 Bounds via $1$-dimensional Extremal $L_{p}$-Embedded Spread
In this section, we will prove Proposition 2.15, which we restate here:
Proposition 6.1 (Restatement of Proposition 2.15).
For a graph $G$ with maximum degree $\Delta$,
$\gamma^{(1)}(G)\lesssim\left[\alpha(G)\right]^{2}\cdot\frac{n^{3}}{\left[\overline{s}_{2}(G)\right]^{2}}.$
When combined with Proposition 2.13 and Proposition 2.14, this will imply the remaining bounds in Theorem 1.5.
To do this, we will use the definition of $L^{p}$-embedded spread, $\overline{Q}_{p}^{(d)}$, given in Definition 4.3. Recall the lemma relating $\overline{Q}_{2}^{(d)}$ and $\gamma^{(d)}$:
Lemma 6.2 (Restatement of Lemma 4.4).
For a graph $G$, $\gamma^{(d)}(G)=\frac{2n}{\overline{Q}^{(d)}_{2}(G)}$.
Given this equivalence, Proposition 2.15 is equivalent to the following lemma:
Lemma 6.3.
Let $G$ be a graph,
$\left[\alpha(G)\right]^{2}\cdot\overline{Q}_{2}^{(1)}(G)\gtrsim\frac{\left[\overline{s}_{2}(G)\right]^{2}}{n^{2}}.$
To prove this, we will apply Cauchy-Schwarz directly to $\overline{s}_{2}(G)$, and then apply a metric embedding to the resulting quantity. The specific metric embedding result we will use was proven by Biswal, Lee, and Rao [BLR10]:
Proposition 6.4 ([BLR10, Theorem 4.4]).
For a metric $(X,d)$ with padded partition modulus $\alpha(X,d)$, there exists an embedding $f:X\to\mathbb{R}$ such that:
  • For every $x,y\in X$, $|f(x)-f(y)|\leq d(x,y)$ (i.e., $f$ is non-expansive).
  • $\alpha(X,d)^{2}\sum_{x,y\in X}|f(x)-f(y)|^{2}\gtrsim\sum_{x,y\in X}d(x,y)^{2}$.
We previously used a more general algorithmic form of this result in the proof of Lemma 4.5.
The first step is given by the following lemma:
Lemma 6.5.
For a graph $G$,
[…]
Proof.
The result follows directly from Cauchy-Schwarz. ∎
The second step is then given by the following lemma:
Lemma 6.6.
For a graph $G$,
[…]